Topic:Acid and Bases

Q1

a) calculate the pH of the following solutions:

i) 0.002 mol dm-3 KOH (2 marks)

ii) 0.10 mol dm-3 C6H5COOH ( K of benzoin acid = 6.3 x 10-5 mol dm-3 ) ( 4 marks)


b) The pH of a 0.10 mol dm-3 solution of a weak mono proton acid, HA , is 2.85. Determine the K of the acid. ( 2 marks)


c) A 500 cm3 solution containing 1.8g of a weak acid HA has a pH of 3.5. Calculate the molar mass of the acid, given that it has a value for K of 2.0 x 10 -6mol dm . ( 5 marks)


please answers all the questions with detailed explanation and steps πŸ™πŸΎπŸ™πŸΎπŸ™πŸΎπŸ™πŸΎπŸ™πŸΎπŸ™πŸΎ


1
Expert's answer
2022-01-19T11:37:01-0500

1 )

a )( i )

0.002 M = 0.002 mol / dm3 KOH Solution


So, Concentration of OH- = 0.002 = 2 Γ— 10-3 .


So, POH = -log 2Γ— 10-3 = 3 - log 2 = 2.7


So , PH = 14 - 2.7 = 11.3


( a ) ( ii ) H+ = under root C . Ka

= under root 0.1 Γ— 6.3 Γ— 10-5

= under root 6.3 Γ— 10-6 .

= 2.6 Γ— 10-3 .

PH = - log 2.6 Γ— 10-3

PH = 3 - log 2.6 = 2.6


( b ) pH = 2.85 .


(H+)2 = (10-2.85)2 = 0.1 Γ— k


So, k = 10--4.7 = 2 Γ— 10-5 .


( c ) pH = 3.5

( H+ ) 2 =( 10-3.5 )2 = c Γ— k

10-7 = ( 1.8 / M Γ— 0.5 ) Γ— 2 Γ— 10-6 .


M = molar mass of weak acid = 72 g / mol.